OpenStudy (angelwings996):

what are all the real zeros y = (x –12)^3 – 7

OpenStudy (angelwings996):

I have done some of my work already but I am not sure if it is right :/

OpenStudy (anonymous):

\[(x-12)^3=7\] \[x-12=\sqrt[3]{7}\] \[x=12+\sqrt[3]{7}\]

OpenStudy (angelwings996):

That is what I have, but what are the real zeros ?

OpenStudy (anonymous):

that is the real zero

OpenStudy (angelwings996):

That's the answer ?

OpenStudy (anonymous):

there is only one real zero, and that is it

OpenStudy (angelwings996):

Oh okay. Thank you!

OpenStudy (anonymous):

this thing crosses the \(x\) axis in only one place it looks just like \(y=x^3\) translated right 12 units, down 7

OpenStudy (angelwings996):

Yeah, I had the answer all along I guess. I just wasn't sure if this was the actual answer or not